Having problem understanding how to write number as a power of the given base:
16; base 2
-216; base -6
2401; base 7
256; base -4
8/27; base 2/3

Answer by lwsshak3(11628)

Having problem understanding how to write number as a power of the given base:
16; base 2
-216; base -6
2401; base 7
256; base -4
8/27; base 2/3
**
raising aħ base to an even power always come out positive
raising a negative base to an odd power always come out negative
..
16; base 2 =2^4
-216; base -6=(-6)^3
2401; base 7=7^4
256; base -4=(-4)^4
8/27; base 2/3=(2/3)^3